OpenStudy (anonymous):

f(x) = x-5 g(x)= x^3 h(x)= √x^2+1 Write this expession as a composition, using only the functions f,g and h. √ (x^3 - 5)^2 + 1 Does anyone know how to do this please? I have an exam coming up and I would appreciate it.

OpenStudy (amistre64):

the copy and paste technique tends to cut off the top of the radical. Can you tell us what is spossed to be 'under' the radical?

OpenStudy (amistre64):

(h o f o g)(x) is one way to write this thing

OpenStudy (amistre64):

h(f(x)) = sqrt. f(x)^2 + 1 f(g(x)) = g(x) - 5 g(x) = x^3 its like nesting functions; you fit them into each other like a set of russian dolls

OpenStudy (anonymous):

It is all meant to be under it. I was thinking that was the answer but I don't know how to work it out. Could you please tell me how you got the answer?

OpenStudy (anonymous):

Thank you so much for your help. I understand it a lot better now.

OpenStudy (amistre64):

i simply used the facts they gave and tore the new equation apart to investigate its inner workings

OpenStudy (amistre64):

we had to rip out the 'x' part and re constuct it with the missing pieces

OpenStudy (amistre64):

sqrt. ______ ^2 +1 ; h(x) matches the profile , h(f(g(x))) /||||||||\ _______ - 5 ; f(x) matches the profile, f(g(x)) /|||||||||\ ----- ; g(x) matches the profile

OpenStudy (anonymous):

Yeah that makes sense thanks.